### v3

@@ 3 Preliminaries
In this section, we first introduce the notation that is used in both family-driven search (FS) and caseworker-driven search (CS). We then describe the processes of the two different search technologies and derive characterizations of agents’ utilities.
@@ 3.1 Model
In our model, agents (children and families) have observable characteristics. Agents with the same characteristics are said to be of the same type. We treat sibling sets of children who should be placed together as a single child. Furthermore, we see the caseworker as a direct representative of the child, i.e., we think of a child-caseworker pair as one child agent. We let $C=\{c_{1},\ldots,c_{n}\}$ and $F=\{f_{1},\ldots,f_{m}\}$ denote the set of all $n$ child types and the set of all $m$ family types, respectively. Individual agents are indifferent between agents of the same type, i.e., their preferences are over agent types: A child of type $c$ has a value $v_{c}({f})\in\mathbb{R}$ for family type $f$, and a family of type $f$ has a value $v_{f}({c})\in\mathbb{R}$ for child type $c$. Preferences are assumed to be strict, i.e., $v_{c}({f})\neq v_{c}({f^{\prime}})$ if $f\neq f^{\prime}$ and $v_{f}({c})\neq v_{f}({c^{\prime}})$ if $c\neq c^{\prime}$. Agents’ valuations are summarized by a list of vectors $v=(v_{c_{1}},\ldots,v_{c_{n}},v_{f_{1}},\ldots,v_{f_{m}})$, and each agent has a value of 0 for remaining unmatched. Given valuations $v$, we let $\bar{v}$ denote the maximal value of all $v_{c}({f})$ and $v_{f}({c})$.
There are infinitely many discrete time steps. At any time step, there is at most one agent of each type present in the system. Thus, we can use $c\in C$ and $f\in F$ to refer to either individual agents and agent types without ambiguity. We refer to an agent (or agent type) from either set as $i\in A:=C\cup F$. From now on, we will simply say that agent $i\in A$ is active if they are present at the current time step.
@@
There is uncertainty regarding whether a specific child $c$ and a specific family $f$ are compatible: With probability $p\in(0,1)$, $f$ is a suitable match for $c$ and unsuitable with probability $1-p$. Whether a match is suitable or not is determined independently at random for child-family pairs. We refer to $p$ as the match success probability. Parameter $p$ captures the following aspect of adoption markets: When a family shows interest in a child, the family’s decision is based on limited reported information, such as the sex, ethnicity, age of the child, and known disabilities. However, there are many other important characteristics of a child that determine whether the child is actually a good match for the family and whether there is mutual attraction. The same holds for a child (or his caseworker) showing interest in a family. Only if a family $f$ is a suitable match for child $c$ can a match between $c$ and $f$ be formed. If $c$ and $f$ form a match they both obtain a value of $v_{c}({f})$ and $v_{f}({c})$, respectively. Determining the suitability of a match is costly for both sides, as it is a time-consuming process. Children and families incur search costs $\kappa_{C}\in\mathbb{R}_{+}$ and $\kappa_{F}\in\mathbb{R}_{+}$, respectively, each time the suitability of a match including them is determined. All agents discount the future, however, they only discount time steps in which they are active. Children’s and families’ discount factors are $\delta_{C}\in[0,1)$ and $\delta_{F}\in[0,1)$, respectively. [^3] An alternative interpretation is to think of children and families leaving the process before the next time step when they are active with probability $1-\delta_{C}$ and $1-\delta_{F}$, respectively.
@@
In the remainder of this section, we describe the two different search technologies. We analyze the dynamic stochastic games induced by an instance and a search technology in a full information environment.
Family-driven Search (FS): At the beginning of each time step, after a child $c$ is randomly chosen to be active, each family that is active and interested in $c$ lets the caseworker of the child know of their interest. This corresponds to families responding to an email announcement made by a caseworker. The caseworker immediately discards any families without mutual interest in $c$, i.e., where $s_{c}(f)=0$ or $s_{f}(c)=0$. The caseworker then investigates all remaining families to determine whether they would actually be a suitable match. Recall that each investigated family is a suitable match for $c$ with probability $p$—which is determined independently for each family—and that each agent incurs cost for each investigation including them. After all families have been processed by the caseworker, $c$ either matches with the most-preferred choice from all those families that were identified as suitable matches or remains unmatched if no such family exists. We move on to the next time step.
Caseworker-driven Search (CS): After a child $c$ is randomly chosen to be active at the beginning of a time step, the child’s caseworker is presented with a list of active families ordered in decreasing order of $v_{c}({f})$. Families are then sequentially processed. The caseworker skips any family in which $c$ is not mutually interested. If there is mutual interest between $c$ and $f$, the suitability of a match between $c$ and $f$ is investigated. If the match turns out to be suitable, $c$’s search is over, $c$ and $f$ are matched and leave. We then move on to the next time step. Otherwise, the caseworker continues the search by selecting the next family in the list. If all families have been processed, the child remains unmatched and we move on to the next time step. Proposition A.1 in online Appendix A.1 shows that $c$’s utility is maximized if the caseworker processes families in decreasing order of $v_{c}(f)$.
@@ 3.2 Utilities
We define agents’ utilities at a time step and characterize their flow utilities in both FS- and CS-induced stochastic games. Assume child $c$ is active at the current time step and that active families are yet to be determined. Let $f$ be an arbitrary family. For any $s\in S$, let $b_{cf}(s)=|\{f^{\prime}\in M_{c}(s)\mid v_{c}({f^{\prime}})>v_{c}({f})\}|$ denote the number of families in $M_{c}(s)$ that $c$ likes better than $f$. Further, let $\beta_{cf}(s)$ denote the probability that $c$ will not match with any other family that $c$ prefers over $f$ at the current time step. Noting that for any child $c$ the probability that a mutually interested family $f^{\prime}$ is active at the current time step and a suitable match is $\lambda p$, it follows immediately that $\beta_{cf}(s)=(1-\lambda p)^{b_{cf}(s)}$ for both FS and CS. In FS and CS, the expected immediate utility that active child $c$ will obtain at an arbitrary time step is then given by
$\bar{u}^{FS}_{c}(s)=\lambda\sum_{f^{\prime}\in M_{c}(s)}\beta_{cf^{\prime}}(s)pv_{c}({f^{\prime}})-\kappa_{C}$ (1)
and
$\bar{u}^{CS}_{c}(s)=\lambda\sum_{f^{\prime}\in M_{c}(s)}\beta_{cf^{\prime}}(s)\left(pv_{c}({f^{\prime}})-\kappa_{C}\right),$ (2)
respectively. Similarly, for a family $f$, we can express the expected immediate utilities at an arbitrary time step (conditional on $f$ being active) by
$\bar{u}^{FS}_{f}(s)=\frac{1}{n}\sum_{c^{\prime}\in M_{f}(s)}\beta_{c^{\prime}f}(s)pv_{c^{\prime}}({f})-\kappa_{F}$ (3)
for FS and
$\bar{u}^{CS}_{f}(s)=\frac{1}{n}\sum_{c^{\prime}\in M_{f}(s)}\beta_{c^{\prime}f}(s)\left(pv_{c^{\prime}}({f})-\kappa_{F}\right)$ (4)
for CS. From this, the crucial difference between FS and CS in our model becomes apparent: In FS, search costs are always incurred if there is mutual interest between agents. In CS, however, the sequential nature of the search means that search costs are only incurred if there is mutual interest and all previous match attempts at the time step have been unsuccessful.
@@
Given strategy profile $s$, $c$’s utility in FS is the unique value $u^{FS}_{c}(s)$ that satisfies
@@
Similarly, $f$’s utility in FS is the unique value $u^{FS}_{f}(s)$ that satisfies
@@
In CS, $c$’s utility in CS is the unique value $u^{CS}_{c}(s)$ that satisfies
@@
Similarly, $f$’s utility in CS is the unique value $u^{CS}_{f}(s)$ that satisfies
@@
A formal proof can be found in online Appendix B.1. One difference between FS and CS is immediately apparent from the above balance equations. In CS, search costs only incur if previous match attempts have been unsuccessful. In FS, however, costs incur with certainty if there is mutual interest and the family is active.
@@ 4.1 Threshold Strategies
In a CS-TS or an FS-TS, threshold $y_{i}$ can be interpreted as $i$’s reservation utility. [^5] Note that our threshold strategies do not correspond to standard simple threshold strategy (where an agent is willing to match with another agent if and only if their match value is above a certain threshold). This is because we find that agents cannot always maximize their utility by playing simple threshold strategies, as Proposition A.3 in online Appendix A.2 illustrates. Informally this follows because families’ search costs in FS can outweigh the expected benefit from being matched with a very desirable child because of the competition from other families that this child prefers. Let $u^{FS*}_{i}(s_{-i})$ and $u^{CS*}_{i}(s_{-i})$ denote $i$’s utility from a best response to $s_{-i}$ in FS and CS, respectively. As before, we simply write $u^{*}_{i}(s_{-i})$ if there is no ambiguity. Proposition 4.2 shows that agents’ best responses always have the form of a threshold strategy.
@@
A formal proof can be found in online Appendix B.2. To derive results later on, it will prove useful to switch between thresholds and strategies. We therefore provide the following definition.
@@
We again omit the superscript if this does not lead to ambiguity. It is trivial to obtain $s^{CS}(y)$ by inserting $y$ in the corresponding equations in Definition 4.1. Algorithm 1 from online Appendix C can be used to compute $s^{FS}(y)$. From now on, we use $\beta^{FS}_{cf}(y)$ and $\beta^{CS}_{cf}(y)$ as shorthand-notation for $\beta_{cf}(s^{FS}(y))$ and $\beta_{cf}(s^{CS}(y))$, respectively. Whenever it is clear from the context, we simply write $\beta_{cf}(y)$.
@@ 4.2 Equilibrium Existence and Lattice Structure
In this section, we show that Nash equilibria always exist under both search technologies. [^6] Technically, we have a stochastic game model, and therefore, these are Nash equilibria of stochastic games. They correspond to the Markov-perfect Nash-equilibrium selection among subgame-perfect Nash equilibria if we consider the problem as a repeated game. We say that strategy profile $s$ is an equilibrium in FS (FSE) if $s$ is a Nash equilibrium in the game induced by FS. Analogously, strategy profile $s$ is an equilibrium in CS (CSE) if $s$ is a Nash equilibrium in the game induced by CS. We use $S^{FS}$ to denote the set of FSE, and let $Y^{FS}=\{(u_{i}(s))_{i\in A}\}_{s\in S^{FS}}$ be the corresponding set of equilibrium threshold profiles in FS. For CS, those sets are defined analogously. Before we can prove that these sets are never empty, we need to define a partial order $\leq_{C}$ on $Y=[0,\bar{v}]^{n+m}$. Note that if agents only play individually rational strategies their utility is always lower bounded by 00 and upper bounded by $\bar{v}$.
@@
A formal proof can be found in online Appendix B.3. The proof proceeds by defining a best-response mapping and showing fixed-point, and therefore equilibrium, existence using Tarski’s fixed point theorem.
@@ 5.1 Pareto Comparison
A formal proof can be found in online Appendix B.1. Intuitively, if two agents are not mutually interested in each other in CS but are in FS, then at least one of them had to lower their interest threshold in FS—which means that their optimal reservation utility is lower in FS. However, since the optimal reservation utility corresponds to the once-discounted utility, this agent must be strictly worse off.
@@
Proof 5.4
It is quite intuitive that a child $c$ cannot be worse off under CS if all families that are mutually interested in $c$ under FS are also interested in $c$ under CS. This allows us to finally show the following theorem.
@@
A formal proof can be found in online Appendix B.2. The main intuition for why an FSE can never be a Pareto improvement over a CSE is that the only way an agent can be better off in FS compared to CS is to have a higher chance of matching with someone they like. But with Lemma 5.2, this implies that some other agent had to lower their interest threshold, which means that their utility decreased. There are two reasons why a CSE can Pareto dominate an FSE. First, CS can save agents search costs. Second, because search costs are only incurred in CS if previous match attempts at the current time step have been unsuccessful, agents do not have to worry about accumulating search costs that much in CS unlike in FS. Therefore, agents are incentivized to express interest in more potential match candidates in CS compared to FS.

### v4

@@ 4 Preliminaries
In this section, we develop an analytical model to contrast the two search disciplines and derive characterizations of agents’ utilities.
@@ 4.1 Model
In our model, agents (children and families) have observable characteristics. Agents with the same characteristics are said to be of the same type. We treat sibling sets of children who should be placed together as a single child. Furthermore, we view the caseworker as a direct representative of the child; thus, we consider a child-caseworker pair as a single child agent. We let $C=\{c_{1},\ldots,c_{n}\}$ and $F=\{f_{1},\ldots,f_{m}\}$ denote the set of all $n$ child types and the set of all $m$ family types, respectively. Individual agents are indifferent between agents of the same type, i.e., their preferences are over agent types: A child of type $c$ has a value $v_{c}({f})\in\mathbb{R}$ for family type $f$, and a family of type $f$ has a value $v_{f}({c})\in\mathbb{R}$ for child type $c$. Preferences are assumed to be strict, i.e., $v_{c}({f})\neq v_{c}({f^{\prime}})$ if $f\neq f^{\prime}$ and $v_{f}({c})\neq v_{f}({c^{\prime}})$ if $c\neq c^{\prime}$. Agents’ valuations are summarized by a list of vectors $v=(v_{c_{1}},\ldots,v_{c_{n}},v_{f_{1}},\ldots,v_{f_{m}})$, and each agent has a value of 0 for remaining unmatched. Given valuations $v$, we let $\bar{v}$ denote the maximum value of all $v_{c}({f})$ and $v_{f}({c})$.
There are infinitely many discrete time steps. At any time step, there is at most one agent of each type present in the system. Thus, we can use $c\in C$ and $f\in F$ to refer to either individual agents and agent types without ambiguity. We refer to an agent (or agent type) from either set as $i\in A:=C\cup F$. From now on, we will simply say that agent $i\in A$ is active if they are present at the current time step. To account for multiple agents from the same type being present, we can introduce multiple agent types that are arbitrarily close in value, effectively creating a tie-breaker between agents of “almost” the same type.
@@
There is uncertainty regarding whether a specific child $c$ and a specific family $f$ are compatible: With probability $p\in(0,1)$, $f$ is a suitable match for $c$ and unsuitable with probability $1-p$. Whether a match is suitable or not is determined independently at random for child-family pairs. We refer to $p$ as the match success probability. Parameter $p$ captures the following aspect of adoption markets: When a family shows interest in a child, the family’s decision is based on limited reported information, such as the sex, ethnicity, age of the child, and known disabilities. However, there are many other important characteristics of a child that determine whether the child is actually a good match for the family and whether there is mutual attraction. The same holds for a child (or his caseworker) showing interest in a family. Only if a family $f$ is a suitable match for child $c$ can a match between $c$ and $f$ be formed. If $c$ and $f$ form a match, they both obtain a value of $v_{c}({f})$ and $v_{f}({c})$, respectively. [^3] Note that a uniform, binary suitability probability $p$ is an abstraction made to keep the model tractable. See online Appendix B. Determining the suitability of a match is costly for both sides, as it is a time-consuming process. Children and families incur search costs $\kappa_{C}\in\mathbb{R}_{+}$ and $\kappa_{F}\in\mathbb{R}_{+}$, respectively, each time the suitability of a match including them is determined. All agents discount the future; however, they only discount time steps in which they are active. Children’s and families’ discount factors are $\delta_{C}\in[0,1)$ and $\delta_{F}\in[0,1)$, respectively. [^4] An alternative interpretation is to think of children and families leaving the process before the next time step when they are active with probability $1-\delta_{C}$ and $1-\delta_{F}$, respectively. Although we keep homogeneous costs and discount factors to keep the analysis and notation concise, they can be made type-specific without loss of generality, and all our results go through.
@@
In the remainder of this section, we describe the two search technologies. We analyze the dynamic stochastic games induced by an instance and a search technology in a full information environment.
Family-driven Search (FS): At the beginning of each time step, after a child $c$ is randomly chosen to be active, each family that is active and interested in $c$ lets the child’s caseworker know of their interest. This corresponds to families responding to an email announcement made by a caseworker. The caseworker immediately discards any families without mutual interest in $c$, i.e., where $s_{c}(f)=0$ or $s_{f}(c)=0$. The caseworker then investigates all remaining families to determine whether they would actually be a suitable match. Recall that each investigated family is a suitable match for $c$ with probability $p$—which is determined independently for each family—and that each agent incurs cost for each investigation in which they are involved. After all families have been processed by the caseworker, $c$ either matches with the most preferred choice from among those families identified as suitable matches or remains unmatched if no such family exists. We move on to the next time step.
Caseworker-driven Search (CS): After a child $c$ is randomly chosen to be active at the beginning of a time step, the child’s caseworker can sequentially inform any of the active families. We assume families are informed in decreasing order of $v_{c}({f})$. [^5] While this is trivially optimal in our model, higher degrees of uncertainty may warrant contacting families in a different order. See online Appendix B for a discussion. If there is mutual interest between $c$ and $f$, the suitability of a match between $c$ and $f$ is investigated. If the match turns out to be suitable, $c$’s search is over, $c$ and $f$ are matched and leave. [^6] In practice, both sides still need to agree to the match. However, given that it would not be rational to investigate a match you are not willing to accept and that the remaining families all have lower values, refusing a suitable match at this point is never rational. We then move on to the next time step. Otherwise, the caseworker continues the search by selecting the next family in the list. If all families have been processed, the child remains unmatched and we move on to the next time step. Proposition C.1 in online Appendix C.1 shows that $c$’s utility is maximized if the caseworker processes families in decreasing order of $v_{c}(f)$.
@@ 4.2 Utilities
We define agents’ utilities at a time step and characterize their flow utilities in both FS- and CS-induced stochastic games. Assume child $c$ is active at the current time step and that active families are yet to be determined. Let $f$ be an arbitrary family. For any $s\in S$, let $b_{cf}(s)=|\{f^{\prime}\in M_{c}(s)\mid v_{c}({f^{\prime}})>v_{c}({f})\}|$ denote the number of families in $M_{c}(s)$ that $c$ likes better than $f$. Further, let $\beta_{cf}(s)$ denote the probability that $c$ will not match with any other family that $c$ prefers over $f$ at the current time step. Noting that for any child $c$ the probability that a mutually interested family $f^{\prime}$ is active at the current time step and a suitable match is $\lambda p$, it follows immediately that $\beta_{cf}(s)=(1-\lambda p)^{b_{cf}(s)}$ for both FS and CS.
In FS, since investigations are conducted simultaneously, search costs are incurred for all of them, and therefore, the expected immediate utility for the active child $c$ at an arbitrary time step follows as
$\bar{u}^{FS}_{c}(s)=\lambda\sum_{f^{\prime}\in M_{c}(s)}\Big{(}\beta_{cf^{\prime}}(s)pv_{c}({f^{\prime}})-\kappa_{C}\Big{)}.$ (1)
In CS, as investigations are conducted sequentially, search costs $\kappa_{C}$ are only incurred if the child has not successfully matched with a higher-valued family. Consequently, the expected immediate utility of the active child $c$ follows as
$\bar{u}^{CS}_{c}(s)=\lambda\sum_{f^{\prime}\in M_{c}(s)}\beta_{cf^{\prime}}(s)\left(pv_{c}({f^{\prime}})-\kappa_{C}\right).$ (2)
In both cases above, $\lambda$ is the probability that a family is present to be investigated when the child is active.
Similarly, for a family $f$, we can express the expected immediate utilities for FS and CS at an arbitrary time step (conditional on $f$ being active) by
$\bar{u}^{FS}_{f}(s)=\frac{1}{n}\sum_{c^{\prime}\in M_{f}(s)}\Big{(}\beta_{c^{\prime}f}(s)pv_{c^{\prime}}({f})-\kappa_{F}\Big{)}$ (3)
$\bar{u}^{CS}_{f}(s)=\frac{1}{n}\sum_{c^{\prime}\in M_{f}(s)}\beta_{c^{\prime}f}(s)\left(pv_{c^{\prime}}({f})-\kappa_{F}\right),$ (4)
where $1/n$ is the probability that a given child is active in a timestep.
From this, the crucial difference between FS and CS in our model becomes apparent: In FS, search costs are always incurred if there is mutual interest between agents. In CS, however, the sequential nature of the search means that search costs are only incurred if there is mutual interest and all previous match attempts at the time step have been unsuccessful.
@@
Given strategy profile $s$, child $c$’s utility in FS is the unique value $u^{FS}_{c}(s)$ that satisfies
@@
Similarly, family $f$’s utility in FS is the unique value $u^{FS}_{f}(s)$ that satisfies
@@
In CS, child $c$’s utility in CS is the unique value $u^{CS}_{c}(s)$ that satisfies
@@
Similarly, family $f$’s utility in CS is the unique value $u^{CS}_{f}(s)$ that satisfies
@@
A formal proof can be found in online Appendix D.1. One difference between FS and CS is immediately apparent from the above balance equations. In CS, search costs only incur if previous match attempts have been unsuccessful. In FS, however, costs are always incurred if a family is present and mutually interested.
@@ 5.1 Threshold Strategies
In a CS-TS or an FS-TS, the threshold $y_{i}$ can be interpreted as $i$’s reservation utility. However, note that these threshold strategies are more involved than standard simple threshold strategies (Adachi 2003, Immorlica et al. 2023). While in a simple threshold strategy, an agent would be interested if the expected value is above their reserve utility, i.e., $pv_{i}({j})\geq y_{i}$, this does not suffice in this case. Instead, agents also have to account for their costs and for the likelihood that the child may find another match with a higher value during the same period. [^8] See Proposition C.3 in online Appendix C.2 for an illustration of the non-existence of simple threshold best responses.
Let $u^{FS*}_{i}(s_{-i})$ and $u^{CS*}_{i}(s_{-i})$ denote $i$’s utility from a best response to $s_{-i}$ in FS and CS, respectively. As before, we simply write $u^{*}_{i}(s_{-i})$ if there is no ambiguity. Proposition 5.2 shows that agents’ best responses always have the form of a threshold strategy.
@@
A formal proof can be found in online Appendix D.2. To derive results later on, it will prove useful to switch between thresholds and strategies. We therefore provide the following definition.
@@
We again omit the superscript if this does not lead to ambiguity. It is trivial to obtain $s^{CS}(y)$ by inserting $y$ in the corresponding equations in Definition 5.1. Algorithm 1 from online Appendix E can be used to compute $s^{FS}(y)$. From now on, we use $\beta^{FS}_{cf}(y)$ and $\beta^{CS}_{cf}(y)$ as shorthand-notation for $\beta_{cf}(s^{FS}(y))$ and $\beta_{cf}(s^{CS}(y))$, respectively. Whenever it is clear from the context, we simply write $\beta_{cf}(y)$.
@@ 5.2 Equilibrium Existence and Lattice Structure
In this section, we show that Nash equilibria always exist under both search technologies. [^9] Technically, we have a stochastic game model, and therefore, these are Nash equilibria of stochastic games. They correspond to the Markov-perfect Nash-equilibrium selection among subgame-perfect Nash equilibria when considered as a repeated game. We say that strategy profile $s$ is an equilibrium in FS (FSE) if $s$ is a Nash equilibrium in the game induced by FS. Analogously, strategy profile $s$ is an equilibrium in CS (CSE) if $s$ is a Nash equilibrium in the game induced by CS. We use $S^{FS}$ to denote the set of FSE, and let $Y^{FS}=\{(u_{i}(s))_{i\in A}\}_{s\in S^{FS}}$ be the corresponding set of equilibrium threshold profiles in FS. For CS, those sets are defined analogously. Before we can prove that these sets are never empty, we need to define a partial order $\leq_{C}$ on $Y=[0,\bar{v}]^{n+m}$. Note that if agents only play individually rational strategies, their utility is always lower bounded by 00 and upper bounded by $\bar{v}$.
@@
A formal proof can be found in online Appendix D.3. The proof proceeds by defining a best-response mapping and showing fixed-point existence using Tarski’s fixed-point theorem (Tarski 1955). Such fixed points coincide with pure-strategy equilibria.
@@ 6.1 Pareto Comparison
A formal proof can be found in online Appendix D.1. Intuitively, if two agents are not mutually interested in each other in CS but are in FS, then at least one of them had to lower their interest threshold in FS—which means that their optimal reservation utility is lower in FS. However, since the optimal reservation utility corresponds to the once-discounted utility, this agent must be strictly worse off.
@@
A formal proof can be found in online Appendix D.2.
It is quite intuitive that a child $c$ cannot be worse off under CS if all families that are mutually interested in $c$ under FS are also interested in $c$ under CS. This allows us to show the following theorem as our main result.
@@
A formal proof can be found in online Appendix D.3. The main intuition for why an FSE can never be a Pareto improvement over a CSE is that the only way an agent can be better off in FS compared to CS is to have a higher chance of matching with someone they like. But with Lemma 6.2, this implies that some other agent had to lower their interest threshold, which means that their utility decreased. There are two reasons why a CSE can Pareto dominate an FSE. First, CS can save agents’ search costs. Second, because search costs are only incurred in CS if previous match attempts at the current time step have been unsuccessful, agents do not have to worry about accumulating search costs as much in CS as in FS. Therefore, agents are incentivized to express interest in more potential match candidates in CS compared to FS.